Juanita has 28,000 to invest in two accounts that pay simple interest on an annual basis. One account pays 5% in simple interest and the other account pays 2%. How much would she have to invest in each account to earn a total interest of $830.00 after one year?
Answer by richwmiller(17219)

x+y=28000,
0.05*x+0.02*y=830
x=28000-y
0.05*(28000-y)+0.02*y=830
1400-0.05y+0.02*y=830
-0.03*y=-570
y= 19000
x=28000-y
x=9000
check
0.05*9000+0.02*19000=830
450+380=830
830=830
ok